OpenStudy (anonymous):

4a^2-36b^2 factored out

OpenStudy (radar):

This is the difference of two perfect squares. \[4a ^{2}-36b ^{2}\]\[(2a-6b)(2a+6b) = 2(a-3b)2(a+3b)=4(a-3b)(a+3b)\]